OpenStudy (anonymous):

Which ordered pair is a solution of the equation y=x-4 ?

OpenStudy (anonymous):

(2,6) (6,2) (-2,6) (-6,2)

OpenStudy (anonymous):

Plug in the points

OpenStudy (anonymous):

Point #1: 6 = 2 - 4 Point #2 2 = 6 - 4 Point #3 6 = -2 - 4 Point #4 2 = -6 - 4 Which one is true?

OpenStudy (nirmalnema):

b option is correct

OpenStudy (anonymous):

B is true because 6 - 4 = 2